OpenStudy (anonymous):

solve the equations x+3=10/x and 6x/x-6 - 4/x = 24/ x^2-6x

OpenStudy (anonymous):

Your top Equation looks like this :\[x + 3 = \frac{ 10 }{ x }\] Knowing that your trying to find what x is (Simplify) Your answer should look like this : x= -5 x=2 Your 2nd Equation looks like this \[6 \times \frac{ x }{ x } - 6 - \frac{ 4 }{ x }=\frac{ 24 }{ x^2 } -6x\] Of course simplify and you would get : \[-\frac{ 4 }{ x }=\frac{ 24 }{ x^2 }-6x\] Hope this helped!

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

No problem (;